# Cross-Lingual Transfer of Debiasing and Detoxification in Multilingual LLMs: An Extensive Investigation

### v2

@@ 5 Additional Analyses
Prior work has shown that the extent to which models exhibit biased behavior differs significantly across bias categories (Neplenbroek et al., 2024), and that biases included in the finetuning dataset are more effectively mitigated than those that are not (Prakash and Lee, 2023). In this section we explore whether we can replicate the latter finding, and break down the bias scores by bias category for MBBQ and StereoSet, which together cover the bias types in the bias mitigation datasets. In Figure 4 we display the change in bias scores on the MBBQ dataset separated by type of bias, language and finetuning method for Llama 3.1 Instruct, as this model benefited most from debiasing. In Appendix F we separate the change in bias scores for Llama 3.1 Instruct evaluated on StereoSet per bias type (Figure 8). We see that supervised finetuning on the Panda dataset decreases bias scores for most bias types. In particular, we note that strong debiasing takes place for age and gender bias, two of the three types of bias included in the Panda dataset. This is particularly visible in English, but the strong debiasing for age bias also transfers to Dutch and Spanish. Whereas the debiasing caused by supervised finetuning on Panda is limited to age and gender bias in English, most other biases also benefit from debiasing in the other languages. Finetuning on the Panda dataset is not as effective for mitigating socio-economic status bias. While overall BiasDPO is less successful at debiasing than Panda, in English and Spanish we observe some debiasing of gender bias, one of the three bias types included in BiasDPO.
@@
In Figure 5 we separate the debiasing results evaluated on CrowS-Pairs (Figure 5(a)) and the detoxification results (Figure 5(b)) per language. [^10] In Figure 9 in Appendix G we show the same figures for StereoSet (Figure 9(a)) and MBBQ (Figure 9(b)). Debiasing and detoxification do not transfer equally well to all languages. In line with findings by Li et al. (2024), transfer is higher for Indo-European languages that share Latin script, vocabulary, and several typological and linguistic features with English, such as Romance languages French and Portuguese, and Germanic languages German and Swedish. The low detoxification performance for Dutch and debiasing performance for Spanish show that transfer does not always take place for languages from similar language families. Rather, it seems that transfer is worse for lower resource languages like Dutch, Maltese, and Catalan, for which models are likely to have seen fewer training data.
@@
To investigate which features of the evaluation language predict whether bias and toxicity mitigation transfer from English, we compute the correlation between the transfer and a variety of features that prior work has found influential for cross-lingual sharing. We include subword overlap with English, an important factor for cross-lingual factual consistency (Qi et al., 2023), and the similarity to English in terms of language family, geography, and syntax, since cross-lingual sharing also occurs for languages that use a different script and therefore have zero subword overlap with English (Pires et al., 2019; Artetxe et al., 2020; Muller et al., 2021; Pfeiffer et al., 2021). Li et al. (2024) find that the cross-lingual success of DPO for toxicity mitigation is high when English and the evaluation language have similar semantic representations, an alignment they speculate is driven by shared linguistic features and sufficient training data. Therefore, we also incorporate bilingual sentence similarity to English, and percentage of language data in the Common Crawl [^11] https://commoncrawl.org/ as features. We include this analysis in Appendix I. The results show a moderate correlation with amount of language data in Common Crawl ($-0.6<r<-0.4$) for all except the Aya models, and with subword overlap and bilingual sentence similarity only for some models, but are otherwise inconclusive.
@@
(a) Mean absolute change in CrowS-Pairs bias score for models finetuned on Panda (vanilla finetuning).
@@
(b) Mean absolute change in expected maximum toxicity (EMT) for models finetuned on DetoxDPO (DPO).
@@
Figure 5: Mean absolute change in bias and toxicity across models, per language. Errorbars indicate 95% confidence intervals.
@@ 6 Conclusion
We finetuned a wide range of models for bias and toxicity mitigation, comparing the effects of English SFT and DPO on the model’s harmfulness and language generation abilities in non-English evaluation languages. Our results show that SFT is more effective for bias mitigation and DPO for toxicity mitigation, at least when evaluated in the manner typically used to assess bias and toxicity. Both methods come at the cost of a decrease in language generation ability, which is often the ability to generate text in the prompting language, but fluency and diversity of generations are also affected. We find that the bias types that are mitigated most align with those that are present in the finetuning datasets, and that differences in mitigation across languages are best explained by differences in the amount of pretraining data per language. Based on our findings, we recommend executing bias and toxicity mitigation in the evaluation language rather than relying on transfer from English, particularly for languages that are underrepresented in pretraining data. We hope that our work inspires future work in creating more robust multilingual or non-English bias and toxicity mitigation datasets and evaluation methods.
@@
For measuring bias and toxicity we were limited to the languages for which benchmark datasets exist. One of the datasets we evaluate on, StereoSet, was (semi-)automatically translated, which may have introduced errors. For evaluating toxicity, we are additionally limited by Perspective API and the languages in which it is available. Perspective API is a black box model that is also subject to its own biases (Sap et al., 2019), though it is frequently retrained to address these, complicating reproducibility (Pozzobon et al., 2023). Due to limited access to GPU resources, we could only train models for a limited number of epochs and sometimes on a subset of the finetuning dataset. We expect the effects we report to be at least as strong when using more data/epochs, as we already observe debiasing and detoxification for 10-30% of our training (see Appendix E). We were also limited by our GPU resources in the evaluation datasets we could include, leading to the omission of PolygloToxicityPrompts that uses long naturally-occurring contexts to elicit toxic generations (Jain et al., 2024).
@@ Appendix B Model descriptions
(Aryabumi et al., 2024) are a family of multilingual LLMs explicitly trained on data in $23$ languages: Arabic, Chinese (simplified & traditional), Czech, Dutch, English, French, German, Greek, Hebrew, Hindi, Indonesian, Italian, Japanese, Korean, Persian, Polish, Portuguese, Romanian, Russian, Spanish, Turkish, Ukrainian, and Vietnamese. The models are finetuned versions of models from the Cohere Command series (Cohere and Cohere for AI, 2024), and have been pretrained and instruction-tuned in those languages on the Aya Collection (Singh et al., 2024). We use the Aya 23 model with 8B parameters.
@@ Appendix C Finetuning details
All models are finetuned using QLORA with rank $64$, alpha $16$ and dropout $0.1$. For Panda and Jigsaw, the supervised finetuning datasets, we used the parameters in Table 3. Finetuning on those datasets takes around $26$ hours on a single NVIDIA A100 GPU. For DPO we have largely adopted the hyperparameters from Li et al. (2024) (Table 4). DPO training on BiasDPO takes around $3$ hours, and DPO training on DetoxDPO takes around $12$ hours, again on a single NVIDIA A100 GPU.
@@
Table 3: Hyperparameters used for Panda and Jigsaw.
@@ Appendix D Initial evaluation
In Table 5 we report the initial diversity and perplexity (before finetuning) of the instruction-tuned models in non-English languages. Aya 23, and Llama 3.1 Instruct generate the least diverse responses with respect to the prompt, which are more fluent than the more diverse generations produced by other models. Mistral 0.3 produces less diverse and less fluent responses. Gemma 2 9B IT generates the most diverse continuations, which are still more fluent than those from Gemma 2 2B IT, and Llama 3 Instruct.
@@
Aya 23 $33.0\pm 14.1$ $22.1\pm 7.7$
Aya Expanse $49.6\pm 21.8$ $46.3\pm 10.9$
Gemma 2 2B IT $58.0\pm 28.7$ $51.4\pm 6.6$
Gemma 2 9B IT $65.5\pm 28.4$ $60.1\pm 4.8$
Llama 3 Instruct $43.1\pm 19.7$ $33.8\pm 7.9$
Llama 3.1 Instruct $29.2\pm 9.4$ $30.5\pm 4.2$
Mistral 0.3 Instruct $27.4\pm 8.8$ $32.4\pm 7.4$
@@
Table 5: Initial perplexity and diversity of instruction-tuned models. Diversity is the percentage of unigrams in the model’s generated continuation that did not occur in the input prompt. The reported scores are averages and standard deviations over all non-English languages.
Appendix E Learning trajectories
@@
Supervised finetuning on the Panda dataset is more effective to mitigate bias, and DPO training on the DetoxDPO dataset to mitigate toxicity, even though these are vastly different training paradigms which affect the model’s weights and the final model’s abilities differently. Moreover, there is a substantial disparity in the sizes of the corresponding datasets. To see how these differences arise during model training, and whether there is already a debiasing or detoxification effect when the model observes significantly less data, we save $10$ checkpoints of Llama 3.1 Instruct during finetuning on each dataset. For these checkpoints we visualize the stereotype score on the StereoSet and CrowS-Pairs dataset, and toxicity, diversity, fluency and language consistency of model outputs in Figure 6. In Figure 7 we show the same figures for Llama 3.1 Instruct finetuned for toxicity mitigation on Jigsaw and DetoxDPO.
@@
In general, for vanilla finetuned models we notice that large changes in bias, toxicity and language generation ability scores occur in the first 20 to 30 percent of bias/toxicity mitigation finetuning. For both models we see an initial sharp decrease in diversity of generations. For Panda this is also visible in the sharp decrease in bias scores and language consistency. Note that the datasets used for supervised finetuning are much larger than those for DPO, so it seems that only a subset of this data is enough to achieve desired debiasing effects. In comparison, DPO trained models undergo more gradual changes that continue throughout the entirety of finetuning. For both DPO trained models we observe a gradual increase in perplexity and an initial decrease in diversity that is later recovered. For non-English languages, diversity even increases beyond its initial value. Comparing the progression of sharp toxicity decrease for DetoxDPO to the much smaller bias decrease achieved by BiasDPO, it does not seem unlikely that a similar debiasing effect could have been achieved if the BiasDPO dataset had been larger.
@@
(c) LM abilities
Panda
(d) LM abilities
BiasDPO
Figure 6: Bias, language consistency, diversity and perplexity of Llama 3.1 Instruct finetuned on Panda (supervised finetuning) and BiasDPO (DPO) for bias mitigation.
@@
(c) LM abilities
Jigsaw
(d) LM abilities
DetoxDPO
Figure 7: Toxicity, diversity, perplexity and language consistency of Llama 3.1 Instruct finetuned on Jigsaw (supervised finetuning) and DetoxDPO (DPO) for toxicity mitigation.
Appendix F Types of bias

### v3

@@ 5 Additional Analyses
Prior work has shown that the extent to which models exhibit biased behavior differs significantly across bias categories (Neplenbroek et al., 2024), and that biases included in the finetuning dataset are more effectively mitigated than those that are not (Prakash and Lee, 2023). In this section we explore whether we can replicate the latter finding, and break down the bias scores by bias category for MBBQ and StereoSet, which together cover the bias types in the bias mitigation datasets. In Figure 4 we display the change in bias scores on the MBBQ dataset separated by type of bias, language and finetuning method for Llama 3.1 Instruct, as this model benefited most from debiasing. In Appendix G we separate the change in bias scores for Llama 3.1 Instruct evaluated on StereoSet per bias type (Figure 8). We see that SFT on the Panda dataset decreases bias scores for most bias types. In particular, we note that strong debiasing takes place for age and gender bias, two of the three types of bias included in the Panda dataset. This is particularly visible in English, but the strong debiasing for age bias also transfers to Dutch and Spanish. Whereas the debiasing caused by SFT on Panda is limited to age and gender bias in English, most other biases also benefit from debiasing in the other languages. SFT on the Panda dataset is not as effective for mitigating socio-economic status bias. While overall BiasDPO is less successful at debiasing than Panda, in English and Spanish we observe some debiasing of gender bias, one of the three bias types included in BiasDPO.
@@
In Figure 5 we separate the debiasing results evaluated on CrowS-Pairs (Figure 5(a)) and the detoxification results (Figure 5(b)) per language. [^10] In Figure 9 in Appendix H we show the same figures for StereoSet (Figure 9(a)) and MBBQ (Figure 9(b)). Debiasing and detoxification do not transfer equally well to all languages. In line with findings by Li et al. (2024), transfer is higher for Indo-European languages that share Latin script, vocabulary, and several typological and linguistic features with English, such as Romance languages French and Portuguese, and Germanic languages German and Swedish. The low detoxification performance for Dutch and debiasing performance for Spanish show that transfer does not always take place for languages from similar language families. Rather, it seems that transfer is worse for lower resource languages like Dutch, Maltese, and Catalan, for which models are likely to have seen fewer training data.
To investigate which features of the evaluation language predict whether bias and toxicity mitigation transfer from English, we compute correlations between the transfer and a variety of features that prior work has found influential for cross-lingual sharing. First, we include percentage of language data in the Common Crawl [^11] https://commoncrawl.org/ to investigate whether transfer is indeed correlated to the amount of language resources available. We also include subword overlap with English, an important factor for cross-lingual factual consistency (Qi et al., 2023), which we compute by segmenting the Flores-200 (Guzmán et al., 2019; Goyal et al., 2022; Team et al., 2022) corpus with each model’s tokenizer before computing the pairwise overlap between English and each other language.
@@
Next, we include the similarity to English in terms of language family, geography, and syntax, since cross-lingual sharing also occurs for languages that use a different script and therefore have zero subword overlap with English (Pires et al., 2019; Artetxe et al., 2020; Muller et al., 2021; Pfeiffer et al., 2021). We obtain each language’s typological features from the URIEL database using lang2vec (Littell et al., 2017), and compute the cosine similarity between each language’s features and those for English. Li et al. (2024) find that the cross-lingual success of DPO for toxicity mitigation is high when English and the evaluation language have similar semantic representations, an alignment they speculate is driven by shared linguistic features and sufficient training data. Therefore, we follow (Li et al., 2024) and for all evaluation datasets compute the average per-layer cosine similarity between sentence representations for each language and English.
For each model, we compute the Spearman correlation between the possible predictors and the amount of debiasing or detoxification taking place as measured on CrowS-Pairs and RTP-LX, the evaluation datasets with at least $8$ languages. We provide more experimental details including all correlation values (Table 10) in Appendix J.
Overall, we do not observe any significant correlations between the change in bias or toxicity and any of the typological features. On the CrowS-Pairs dataset we observe statistically significant correlations for subword overlap and percentage of language data for only one model each, likely due to the limited number of languages included in this benchmark. For toxicity mitigation, we observe low to moderate correlations with subword overlap and bilingual sentence similarity for Gemma 2 9B IT and its base model. However, we are not able to replicate these findings for any of the other models, and even observe a positive correlation for Aya Expanse. Finally, we find significant modererate correlation with amount of language data in Common Crawl ($-0.6<r<-0.4$) for all except the Aya models. This means that languages with more data in Common Crawl, which serves as a proxy for languages with greater linguistic resources, benefit more from transfer of toxicity mitigation. These correlations are stronger for base models, likely because they have been more recently trained on Common Crawl or similar datasets.
@@
(a) Mean absolute change in CrowS-Pairs bias score for models finetuned on Panda (SFT).
@@
(b) Mean absolute change in toxicity score for models finetuned on DetoxDPO (DPO).
@@
Figure 5: Mean absolute change in bias and toxicity per language, averaged across models. Errorbars indicate 95% confidence intervals.
@@ 6 Conclusion
We finetuned a wide range of models for bias and toxicity mitigation, comparing the effects of English SFT and DPO on the model’s harmfulness and language generation abilities in non-English evaluation languages. We effectively mitigate bias using SFT, but our results show that only DPO effectively mitigates toxicity, at least when evaluated in the manner typically used to assess bias and toxicity. Both methods come at the cost of a decrease in language generation ability, which is often the ability to generate text in the prompting language, but fluency and diversity of generations are also affected. We find that differences in mitigation across languages are best explained by differences in the amount of pretraining data per language. Based on our findings, we recommend executing bias and toxicity mitigation in the evaluation language rather than relying on transfer from English, particularly for lower resource languages that are underrepresented in pretraining data. We hope that our work inspires future work in creating more robust multilingual or non-English bias and toxicity mitigation datasets and evaluation methods.
@@
For measuring bias and toxicity we were limited to the languages for which benchmark datasets exist. One of the datasets we evaluate on, StereoSet, was (semi-)automatically translated, which may have introduced errors. For evaluating toxicity, we are additionally limited by Perspective API and the languages in which it is available. Perspective API is a black box model that is also subject to its own biases (Sap et al., 2019), though it is frequently retrained to address these, complicating reproducibility (Pozzobon et al., 2023). Similarly, we were limited by the number of available datasets that explicitly aim to mitigate bias or toxicity. However, from our results we can conclude that whenever successful transfer of bias or toxicity mitigation occurs, regardless of the dataset or type of finetuning, at least one language generation metric suffers. Therefore, in principle, we also expect those results to hold for other bias/toxicity mitigation datasets.
Due to limited access to GPU resources, we could only train models for a limited number of epochs and sometimes on a subset of the finetuning dataset. We expect the effects we report to be at least as strong when using more data/epochs, as we already observe debiasing and detoxification for 10-30% of our training (see Appendix F). We were also limited by our GPU resources in the evaluation datasets we could include, leading to the omission of PolygloToxicityPrompts that uses long naturally-occurring contexts to elicit toxic generations (Jain et al., 2024).
@@ Appendix A Finetuning datasets
Bias
The SFT dataset we use for bias mitigation is Perturbation Augmentation NLP DAtaset (PANDA) (Qian et al., 2022), a human-annotated dataset consisting of 98K sentence pairs of original and demographically perturbed sentences. The perturbed sentences have been modified along the gender, race/ethnicity, and age axes, perturbing a sentence such as “women like shopping” to “men like shopping”. Finetuning on perturbed sentences in general (Bartl and Leavy, 2024), and on the PANDA dataset specifically has been shown to be an effective debiasing technique for mitigating stereotypical model behavior (Prakash and Lee, 2023; Ranaldi et al., 2024). For DPO finetuning we use BiasDPO (Allam, 2024), a manually designed DPO dataset for mitigating gender, race, and religion biases. The dataset consists of 1.1K prompts with biased and unbiased completions.
Toxicity
The SFT dataset we use for toxicity mitigation is the English Jigsaw Civil Comments dataset. [^12] https://www.kaggle.com/c/jigsaw-unintended-bias-in-toxicity-classification These comments from the Civil Comments platform are human-annotated for toxicity, and we follow Xu et al. (2021) in selecting those with less than 10% toxic annotations as our non-harmful dataset. We select a subset of 98K sentences, matching the size of the PANDA dataset. Even though SFT can only increase the probability of non-toxic text, unlike DPO which simultaneously decreases the probability of toxic text, it has proven an effective method for mitigating toxicity (Gehman et al., 2020). Lee et al. (2024) created a DPO dataset specific for toxicity, which has been shown to mitigate toxicity in languages other than English (Li et al., 2024). The dataset consists of 25K prompts from Wikitext-2 (Merity et al., 2017) with synthetic toxic and non-toxic completions. This is the DPO dataset we use, which we will refer to as DetoxDPO.
@@ Appendix B Model descriptions
(Aryabumi et al., 2024) are a family of multilingual LLMs explicitly trained on data in $23$ languages: Arabic, Chinese (simplified & traditional), Czech, Dutch, English, French, German, Greek, Hebrew, Hindi, Indonesian, Italian, Japanese, Korean, Persian, Polish, Portuguese, Romanian, Russian, Spanish, Turkish, Ukrainian, and Vietnamese. The models are finetuned versions of models from the Cohere Command series (Cohere and Cohere for AI, 2024), and have been pretrained and instruction-tuned in those languages on the Aya Collection (Singh et al., 2024b). We use the Aya 23 model with 8B parameters.
@@ Appendix C Finetuning details
All models are finetuned using QLORA with rank $64$, alpha $16$ and dropout $0.1$. For supervised finetuning we used the parameters in Table 5. Finetuning on Panda and Jigsaw takes around $26$ hours on a single NVIDIA A100 GPU, whereas supervised finetuning on the preferred completions of either DPO dataset takes around $2.5$ hours. For DPO we have largely adopted the hyperparameters from Li et al. (2024) (Table 6). DPO training on BiasDPO takes around $3$ hours, and DPO training on DetoxDPO takes around $12$ hours, again on a single NVIDIA A100 GPU.
@@
Table 5: Hyperparameters used for supervised finetuning. We finetune on BiasDPO for $20$ epochs and on Panda, Jigsaw, and DetoxDPO for $1$ epoch.
@@ Appendix D Initial evaluation
In Table 7 we report the initial diversity, perplexity and question-answering ability (before finetuning) of the instruction-tuned models in non-English languages. Aya 23, and Llama 3.1 Instruct generate the least diverse responses with respect to the prompt, which are more fluent than the more diverse generations produced by other models. Mistral 0.3 produces less diverse and less fluent responses. Gemma 2 9B IT generates the most diverse continuations, which are still more fluent than those from Gemma 2 2B IT, and Llama 3 Instruct. Gemma 2 9B IT and Llama 3.1 Instruct display the best question-answering abilities across languages, compared to Gemma 2 2B IT and Mistral 0.3 Instruct which perform worst on Global-MMLU.
@@
Question-
answering
Aya 23 $33.0\pm 14.1$ $22.1\pm 7.7$ $47.8\pm 2.5$
Aya Expanse $49.6\pm 21.8$ $46.3\pm 10.9$ $54.4\pm 2.8$
Gemma 2 2B IT $58.0\pm 28.7$ $51.4\pm 6.6$ $45.5\pm 3.2$
Gemma 2 9B IT $65.5\pm 28.4$ $60.1\pm 4.8$ $63.4\pm 2.7$
Llama 3 Instruct $43.1\pm 19.7$ $33.8\pm 7.9$ $52.7\pm 4.2$
Llama 3.1 Instruct $29.2\pm 9.4$ $30.5\pm 4.2$ $56.0\pm 3.7$
Mistral 0.3 Instruct $27.4\pm 8.8$ $32.4\pm 7.4$ $46.8\pm 6.6$
@@
Table 7: Initial perplexity, diversity and question-answering ability of instruction-tuned models. Diversity is the percentage of unique unigrams in the model’s generated continuation that did not occur in the input prompt. The reported scores are averages and standard deviations over all non-English languages.
Appendix E Toxicity scores
In Table 8 we present the toxicity scores obtained when including all completions generated by each model next to those obtained when only including the completions in the evaluation language. We only find minimal differences, as the most toxic completions that drive the toxicity scores are those in the prompt language.
Model Finetuning All completions
Only evaluation
language completions
Aya 23 None 0.550 0.541
Aya 23 SFT, Jigsaw 0.589 0.570
Aya 23 SFT, DetoxDPO 0.591 0.576
Aya 23 DPO, DetoxDPO 0.323 0.318
Aya Expanse None 0.480 0.472
Aya Expanse SFT, Jigsaw 0.547 0.535
Aya Expanse SFT, DetoxDPO 0.548 0.534
Aya Expanse DPO, DetoxDPO 0.292 0.284
Gemma 2 9B IT None 0.481 0.495
Gemma 2 9B IT SFT, Jigsaw 0.511 0.559
Gemma 2 9B IT SFT, DetoxDPO 0.593 0.571
Gemma 2 9B IT DPO, DetoxDPO 0.456 0.442
Llama 3.1 Instruct None 0.551 0.539
Llama 3.1 Instruct SFT, Jigsaw 0.590 0.571
Llama 3.1 Instruct SFT, DetoxDPO 0.589 0.566
Llama 3.1 Instruct DPO, DetoxDPO 0.367 0.341
Table 8: Average toxicity scores across languages when including all completions and when only including completions in the evaluation language.
Appendix F Learning trajectories
@@
SFT on the Panda dataset is more effective to mitigate bias, and DPO training on the DetoxDPO dataset to mitigate toxicity, even though these are vastly different training paradigms which affect the model’s weights and the final model’s abilities differently. In particular, the KL constraint in DPO’s loss function prevents the finetuned model’s weights from deviating too far from the original model (Rafailov et al., 2023). In addition, there is a substantial disparity in the sizes of the corresponding datasets. To see how these differences arise during model training, and whether there is already a debiasing or detoxification effect when the model observes significantly less data, we save $10$ checkpoints of Llama 3.1 Instruct during finetuning on each dataset. For these checkpoints we visualize the stereotype score on the StereoSet and CrowS-Pairs datasets, question-answering on Global-MMLU, and toxicity, diversity, fluency and language consistency of model outputs in Figure 6. In Figure 7 we show the same figures for Llama 3.1 Instruct finetuned for toxicity mitigation on Jigsaw and DetoxDPO.
@@
In general, for SFT models we notice that large changes in bias, toxicity and language generation ability scores occur in the first 20 to 30 percent of bias/toxicity mitigation finetuning. For both models we see an initial sharp decrease in diversity of generations, and a small decrease in question-answering ability. For Panda this is also visible in the sharp decrease in bias scores and language consistency. Note that the datasets used for SFT are much larger than those for DPO, so it seems that only a subset of this data is enough to achieve desired debiasing effects. In comparison, DPO trained models undergo more gradual changes that continue throughout the entirety of finetuning. For both DPO trained models we observe a gradual increase in perplexity, a very slight but gradual decrease in question-answering abilities and an initial decrease in diversity that is later recovered. For non-English languages, diversity even increases beyond its initial value. Comparing the progression of sharp toxicity decrease for DetoxDPO to the much smaller bias decrease achieved by BiasDPO, it does not seem unlikely that a similar debiasing effect could have been achieved if the BiasDPO dataset had been larger.
@@
(c) LM abilities Panda
(d) LM abilities BiasDPO
Figure 6: Bias, diversity, perplexity, language consistency and question-answering of Llama 3.1 Instruct finetuned on Panda (SFT) and BiasDPO (DPO) for bias mitigation.
@@
(c) LM abilities Jigsaw
(d) LM abilities DetoxDPO
Figure 7: Toxicity, diversity, perplexity, language consistency and question-answering of Llama 3.1 Instruct finetuned on Jigsaw (SFT) and DetoxDPO (DPO) for toxicity mitigation.
Appendix G Types of bias